OpenStudy (anonymous):

-22 = x - 7 will give medal

OpenStudy (anonymous):

i know its super easy but its like 2am and im about to collapse

OpenStudy (anonymous):

-22+7 =x -15 =x

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

welcum